OpenStudy (anonymous):

List the elements of the function F = {(x,y) | 2x + y = 7} if x {positive integers less than 4} and y {positive integers less than 8}. {( ), ( ), ( )} (Remember to include commas)

OpenStudy (anonymous):

\(x\) a positive integer less than 4, can only be 1, 2, 3 right?

OpenStudy (anonymous):

if \(x=1\) we get \(2\times 1+y=7\implies 2+y=7 \implies y=5\) so one pair is \((1,5)\) try it with \(x=2\) and see what you get for \(y\)